A horizontal vinyl record of mass 0.0592 kg and radius 0.144 m rotates freely about a vertical axis through its center with an angular speed of 4.39 rad/s and a rotational inertia of 6.41 x 10-4 kg·m2. Putty of mass 0.0468 kg drops vertically onto the record from above and sticks to the edge of the record.What is the angular speed of the record immediately afterwards?
1
Expert's answer
2020-06-23T12:56:48-0400

Given, Before putting the Putty of mass "m=0.0468kg" is "\omega_1=4.39rad/s" ,Moment of inertia of Vinyl is "I=6.41\times 10^{-4}kg\cdot m^2" .

Radius of horizontal Vinyl is "r=0.144m" .

Now, After putting the Putty on Vinyl we have Moment of inertia is

"I_2=I+mr^2=16.11\times 10^{-4} \: kg\cdot m^2"

Now, As no external torque is acting thus Angular momentum is conserved,hence,


"I\omega_1=I_2\omega"

Hence,

"\omega=\frac{I}{I_2}\omega_1\implies \omega\approx1.75rad/s"

Thus, required angular velocity is "1.75rad\/s"